40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 289     Entry time: Wed Feb 1 16:58:10 2017
Author: Gabriele 
Type: General 
Category: Characterization 
Subject: Excitation tests 

The excitation is working fine. However, there is a large cross coupling in the excitation between channels: so if I switch on only one of the ESD paths, I actually excite all other disks too.  It might be due to the common ground or to the fact that the cables run close to each other. This needs some investigation, but it's not a big issue for the moment being.

On the other hand, the optical levers are very well decoupled: each one sees a different set of modes. So there is no measurable cross coupling between the disks or the readout.

There seems to be something fishy with the Y picomotor of QPD3: it doesn't always move in the same direction for the same set of steps. Some investigations needed here too. Autocentering might fail, but it's still possible to center it manually.

Excitation at 4:56pm (quiet time before excitation 1170032165, quiet time after excitation 1170032234).

Another excitation at 8:30am (2017-02-02) (quiet time before excitation 1170088185, quiet time after excitation 1170088373)

ELOG V3.1.3-